OpenStudy (anonymous):

wat is 23,000 rouned to the nearset dollor within 1%

OpenStudy (anonymous):

.01 * 23000 = 230 23000 + 230 = 23230 23000 - 230 = 22770

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

you're welcome.